# Model-Based Reward Shaping for Adversarial Inverse Reinforcement Learning in Stochastic Environments

### v1

@@ 5.2 Algorithm Framework
6:        Pretrain transition model $\hat{\mathcal{T}}$.
7:     else
8:        Sample state-action batch from $\mathcal{D}_{exp},\mathcal{D}_{env}$ respectively.
9:        Train $D_{\theta}$ via cross entropy loss  Eq. 9 to classify expert data from samples.
10:        Update dynamic model $\hat{\mathcal{T}}$ with MLE loss and generate $H$-steps trajectories to $\mathcal{D}_{gen}$.
11:        Sample state-action batches from $\mathcal{D}_{env},\textrm{and }\mathcal{D}_{gen}$ with varying ratio.
12:        Update $\pi$ with respect to $\hat{R}_{\theta}$ using Soft Actor Critic policy optimization.
13:     end if
14:  end for
In this section, we present the overall framework of Model-Enhanced Adversarial IRL and illustrate how transition model training is incorporated into the learning loop. We assume the estimated transition distribution $\hat{\mathcal{T}}(\cdot|s,a)$ follows a Gaussian distribution with mean and standard deviation parameterized by the MLP, and the model is updated with standard maximum likelihood loss. At each iteration, the updated transition model is utilized for reward calculation and synthetic data generation, which is stored in the synthetic trajectory replay buffer. Unlike AIRL and GAIL, our framework operates in an off-policy manner, where samples used for both discriminator and policy update are drawn from a combination of the environmental replay buffer and the synthetic replay buffer. An overview of our framework is shown in  Fig. 1, and detailed algorithmic steps are provided in  Alg. 1 and Appendix D.
Sample Efficiency: To improve sample efficiency, we leverage the estimated transition model to generate $H$-steps synthetic trajectories data alongside real interaction data, facilitating policy optimization. Given that the estimated transition model is inaccurate at the beginning, we employ a dynamic ratio between real and synthetic data to prevent the model from being misled by unlikely synthetic transitions (Janner et al., 2019; Zhan et al., 2023). Specifically, early-stage generated trajectories are not stored persistently, unlike real interactions which are fully stored in the off-policy environmental replay buffer. To maintain training stability, we use a synthetic replay buffer with a size that gradually increases as training progresses, ensuring a balanced inclusion of synthetic data over time. The growth rates of the data ratio and buffer size are adjusted based on the complexity of the transition model learning process and can be fine-tuned via hyperparameters.
Distribution Shift: To mitigate distribution shift (Lee et al., 2020; Lin et al., 2020) during training, we employ a strategy involving the learned transition model. Typically, during interaction, the real state $s_{t}$ is used as input to the actor, and the resulting action $a_{t}$ is applied in the environment. To incorporate the transition model, we predict a synthetic state $\hat{s}_{t}$ from previous $s_{t-1}$ and $a_{t-1}$. This generated $\hat{s}_{t}$ is then fed into the actor to produce action $\hat{a}_{t}$. The actions $a_{t}$ and $\hat{a}_{t}$ are mixed and applied to the environment with a certain ratio, and the resulting pairs $(s_{t},a_{t})$ or $(s_{t},\hat{a}_{t})$ are stored in the environmental replay buffer. This approach helps balance the exploration of real and model-predicted dynamics, reducing the impact of distributional discrepancies.
5.3 Performance Analysis
[…]
Assumption 5.2 (Transition Model Error).
Since transition model is trained through a supervised fashion, we can use a PAC generalization bound (Shalev-Shwartz & Ben-David, 2014) for sample error. Therefore, we assume that the total variation distance between $\mathcal{T}$ and $\hat{\mathcal{T}}$ is bounded by $\epsilon_{\mathcal{T}}$ through $[0,T]$:
$\max_{t}\mathbb{E}_{s\sim\pi_{D,t}}\left[D_{TV}(\mathcal{T}(s^{\prime}|s,a)|\hat{\mathcal{T}}(s^{\prime}|s,a))\right]\leq\epsilon_{\mathcal{T}},$ (10)
which is a common assumption that adopted in literature (Janner et al., 2019; Sikchi et al., 2022).
Next, with the assumed total visitation bound on transition models ( 5.2), we can bound the error on rewards learning through model-enhanced reward shaping.
Theorem 5.3 (Reward Function Error Bound).
Proof of  Thm. 5.3 can be found in Appendix C.3. With rewards bound above, we can extend the bound to the value function, which represents the performance difference brought up by estimated transition model error under RL setting.
Theorem 5.4 (Performance Difference Bound).
The performance difference between the optimal policies ($\pi^{*}$ and $\hat{\pi}^{*}$) in corresponding MDPs ($\mathcal{M}^{\prime}\cup R$ and ^$\hat{\mathcal{M}}^{\prime}\cup\hat{R}$) can be bounded as follows:
$\|V_{\mathcal{M}^{\prime}\cup R}^{\pi^{*}}-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\hat{\pi}^{*}}\|\leq\epsilon_{\mathcal{T}}\left[\frac{\gamma}{(1-\gamma)^{2}}R_{\max}+\frac{1+\gamma}{(1-\gamma)^{2}}R_{\max}|\mathcal{S}|\right].$ (12)
The detailed proof of  Thm. 5.4 is presented in Appendix C.6. The above theorem highlights the relationship between the performance gap and the transition model error, also implying that a perfectly-learned transition model ($\epsilon_{\mathcal{T}}\rightarrow 0$) could make the performance difference negligible.
6 Experiments
In this section, we evaluate the performance and sample efficiency of our Model-Enhanced Adversarial IRL framework. We aim to demonstrate the superiority of our method in stochastic environments, achieving better performance and sample efficiency compared to existing approaches. Additionally, in deterministic settings, our method not only maintains competitive performance, but also achieves better sample efficiency with baselines. All experiments are conducted on the MuJoCo benchmarks (Todorov et al., 2012). To simulate stochastic dynamics in MuJoCo, we introduce the agent-unknown Gaussian noise with a mean of 0 and a standard deviation of 0.5 to the environmental interaction steps. All the expert data are generated by an expert trained with standard SAC under deterministic or stochastic MuJoCo environments. Our experiments are designed to highlight the key advantages of our framework:
  • Superiority in Stochastic Environments: In stochastic settings, our method significantly outperforms other approaches, consistently surpassing expert-level performance more rapidly. This enhanced ability to learn under uncertainty is attributed to our framework’s effectiveness in leveraging model-based synthetic data.
  • Performance in Deterministic Environments: We demonstrate that our method is either better or competitive with existing AIL methods’ performances in deterministic settings.
  • Sample Efficiency For both deterministic and stochastic settings, our method can reach expert performance with fewer training steps than all the other baselines.
We primarily compare our approach with other Adversarial Imitation Learning (AIL) methods, including the on-policy algorithms GAIL (Ho & Ermon, 2016) and AIRL (Fu et al., 2017), and the off-policy method Discriminator Actor-Critic (DAC) (Kostrikov et al., 2018). For policy optimization, we use Proximal Policy Optimization (PPO) (Schulman et al., 2017) for both GAIL and AIRL, and Soft Actor-Critic (SAC) (Haarnoja et al., 2018) for both DAC and our proposed method. All implementations of PPO and SAC are referenced from the Clean RL library (Huang et al., 2022). Each algorithm is trained with 100k environmental steps and evaluated each 1k steps across 5 different seeds for tasks including InvertedPendulum-v4 and InvertedDoublePendulum-v4. For Hopper-v3, AIRL and GAIL are trained with 10M steps and evaluated each 100k steps across 5 different seeds, but DAC and our algorithm are trained with 1M environmental steps and evaluated each 10k steps across 5 different seeds. We conduct the experiments under 3 different random seeds to seek more general trend and conclude systematic results. All the experiments are run on the Desktop equipped with RTX 4090 and Core-i9 13900K. The learning curves of all methods are provided in Appendix E.
Performance in Stochastic MuJoCo.
In  Table 2, we present the performance under the stochastic MuJoCo environments. Our method can achieve the best performance compared to all of the baselines in stochastic environments. Specifically, for InvertedDoublePendulum-v4, introducing stochasticity into the dynamics significantly degrades the performance. Our method still can still achieve performance better than all baselines. Furthermore, in the stochastic settings, the performance of DAC is decreased significantly, as DAC’s ineffective reward formulation on state-action pairs under uncertain conditions, also resulting a trend of learning instability (Appendix E). For Hopper-v4, our method remarkably surpasses the all baselines with significantly large margins. These results indicate that our approach can effectively learn the better reward function from demonstrations in stochastic environments, resulting in significant performance improvement.
Table 2: Performance of expert and all algorithms in stochastic MuJoCo Environments.
Environment Expert GAIL AIRL DAC Ours
InvertedPendulum-v4(Stoch) $1000.0_{\pm 0.0}$ $1000.0_{\pm 80.8}$ $1000.0_{\pm 25.4}$ $1000.0_{\pm 69.2}$ $1000.0_{\pm 31.2}$
InvertedDoublePendulum-v4(Stoch) $210.1_{\pm 98.6}$ $276.9_{\pm 25.1}$ $234.9_{\pm 38.1}$ $142.3_{\pm 31.2}$ $286.7_{\pm 39.2}$
Hopper-v3(Stoch) $2257.8_{\pm 317.4}$ $2832.0_{\pm 105.5}$ $2834.0_{\pm 333.2}$ $2408.0_{\pm 78.6}$ $3139.0_{\pm 109.9}$
Performance in Deterministic MuJoCo.
The performance of deterministic MuJoCo environments can be found in  Table 3. For the tasks with deterministic dynamics, our method can achieve the performance aligning with all of baselines and the expert in InvertedDoublePendulum-v4, InvertedDoublePendulum-v4 and Hopper-v3. Additionally, our method shows superiority in learning stability (Appendix E) and sample efficiency (Table 4). Therefore, our method shows competitive performance with the baselines in the deterministic environments with better sample efficiency and learning stability.
Table 3: Performance of expert and all algorithms in deterministic MuJoCo Environments.
Environment Expert GAIL AIRL DAC Ours
InvertedPendulum-v4 $1000.0_{\pm 0.0}$ $1000.0_{\pm 0.0}$ $1000.0_{\pm 0.0}$ $1000.0_{\pm 0.0}$ $1000.0_{\pm 0.0}$
InvertedDoublePendulum-v4 $9320.0_{\pm 578.8}$ $9322.0_{\pm 39.6}$ $9359.0_{\pm 88.3}$ $9360.0_{\pm 12.3}$ $9361.0_{\pm 0.2}$
Hopper-v3 $3329.3_{\pm 306.6}$ $3485.0_{\pm 5.6}$ $3472.0_{\pm 3.6}$ $3495.0_{\pm 4.5}$ $3463.0_{\pm 2.7}$
Sample Efficiency.
In  Table 4, we report the number of training steps required to achieve expert-level performance. Our method shows significant superiority in sample efficiency across all of the benchmarks under both deterministic and stochastic settings. Specifically, for InvertedPendulum-v4, all four methods are able to reach expert-level performance. However, our method reaches expert performance in the fewest training steps, $9.7k$ and $11.7k$ for deterministic and stochastic settings respectively. In the case of InvertedDoublePendulum-v4, introducing stochasticity into the dynamics makes it challenging for all algorithms to reach expert-level performance from noisy demonstrations. Especially, the strong baseline, DAC, failed to reach the expert’s performance, while our method can arrive at the expert-level performance within $27.3k$ steps. For Hopper-v3, all four methods achieve expert-level performance, but our method reaches it with the fewest samples whether in deterministic ($195.3k$) or stochastic ($170.0k$) settings.
Table 4: Averaged steps used for all algorithms to reach experts’ performances in deterministic and stochastic Environments. $X$ stands for that the algorithm does not reach the expert performance within given training steps.
Environment GAIL AIRL DAC Ours
InvertedPendulum-v4 $37.7k$ $38.0k$ $15.3k$ $9.7k$
InvertedPendulum-v4(Stoch) $54.0k$ $58.0k$ $11.7k$ $11.7k$
InvertedDoublePendulum-v4 $87.0k$ $91.5k$ $28.3k$ $11.3k$
InvertedDoublePendulum-v4(Stoch) $72.3k$ $75.3k$ X $27.3k$
Hopper-v3 $5400.0k$ $6400.0k$ $243.5k$ $195.3k$
Hopper-v3(Stoch) $4170.0k$ $3200.0k$ $786.7k$ $170.0k$
7 Conclusion
In this paper, we presented a novel model-enhanced adversarial inverse reinforcement learning framework by incorporating model-based techniques with reward shaping, specifically designed to enhance performance in stochastic environments and maintain competitive performance in deterministic setting while significantly improve performance under the traditional AIL approaches. The theoretical analysis provides guarantees on the optimal policy invariance under the transition model involved reward shaping and highlight the relationship between performance gap and transition model error, showing that the gaps becomes negligible with a well-learned model. Empirical evaluations on Mujoco benchmark environments validate the effectiveness of our method, showcasing its superior performance and sample efficiency across different tasks. Future works will focus on further refining the model estimation process to handle more complex and dynamic environments and exploring extensions of the framework to multi-agent and hierarchical reinforcement learning scenarios. Additionally, it would be valuable to investigate the generalization ability of our framework in a transfer learning tasks. Overall, our approach offers a promising direction for advancing model-based adversarial IRL, with the potential to scale to a broader range of real-world applications.
8 Reproducible Statement
This work uses the open-source MuJoCo (Todorov et al., 2012) as the benchmark. The practical implementation of our method is built on the CleanRL repository (Huang et al., 2022). All hyperparameters to reproduce our experimental results, including learning rates and transition model settings, are explicitly listed in Appendix D. For every reported result, we averaged the performance over three random seeds, and the seed initialization is included for exact reproducibility.
Appendix A Reward Shaping Soft Optimal Policy
Theorem A.1.
@@ Appendix A Reward Shaping Soft Optimal Policy
Proof.
According to Soft VI ( Eq. 2), we can expand the representation of $Q^{soft}_{\hat{R}}(s_{t},a_{t})$ as follows.
@@
From above induction, we can tell that $Q^{soft}_{\hat{R}}(s_{t},a_{t})+\phi(s_{t})$ satisfy the soft bellmen update with original $R$. Thus, with simple induction, we can arrive that $Q_{R}^{soft}(s_{t},a_{t})=Q_{\hat{R}}^{soft}(s_{t},a_{t})+\phi(s_{t})$. Then, we can derive the advantage function
@@
Appendix B Adversarial Reward Learning
Proposition B.1.
Consider an undiscounted MDP. Suppose $f_{\theta}$ and $\pi$ at current iteration are the soft-optimal advantage function and policy for reward function $R_{\theta}$. Minimising the cross-entropy loss of the discriminator under generator $\pi$ is equivalent to maximising the log-likelihood under MCE IRL.
Proof.
@@ Appendix B Adversarial Reward Learning
$=\mathbb{E}_{\mathcal{D}_{exp}}\left[\sum_{t=0}^{T-1}\log\pi(a_{t}|s_{t})+\log\rho_{0}+\sum_{t=1}^{T}\mathcal{T}(s_{t+1}|s_{t},a_{t})\right]$
@@
Breaking down above equations with soft VI (Eq. 2), we can arrive the following.
@@
The transitions in the demonstration $\mathcal{D}_{exp}$ will follow $\mathcal{T}$. Thus we can drop the $\mathbb{E}_{\mathcal{T}}$, and $B$ term will cancel out, ending up to 00. To summarize, the gradient of log MLE loss of MCE IRL is the following.
@@
$\log D_{\theta}(s,a)$ $=f_{\theta}(s,a)-\log(\exp\{f_{\theta}(s,a)\}+\pi(a|s)),$
$\log(1-D_{\theta}(s,a))$ $=\log\pi(a|s)-\log(\exp\{f_{\theta}(s,a)\}+\pi(a|s)).$
@@
$\nabla_{\theta}\log D_{\theta}(s,a)$ $=\nabla_{\theta}f_{\theta}(s,a)-\frac{\exp\{f_{\theta}(s,a)\}\nabla_{\theta}f_{\theta}(s,a)}{\exp\{f_{\theta}(s,a)\}+\pi(a|s)},$
$\nabla_{\theta}\log(1-D_{\theta}(s,a))$ $=-\frac{\exp\{f_{\theta}(s,a)\}\nabla_{\theta}f_{\theta}(s,a)}{\exp\{f_{\theta}(s,a)\}+\pi(a|s)}.$
Since $\pi$ is trained by using $f_{\theta}$ as shaped reward, from soft VI we can derive that $\pi^{*}_{f_{\theta}}(a|s)=\exp A^{soft}_{f_{\theta}}(s,a)$. By assumption, we assume that $f_{\theta}$ is the advantage function of $R_{\theta}$, $f_{\theta}(s,a)=A^{soft}_{R_{\theta}}(s,a)$. From Thm 4.2, we know that $A^{soft}_{R_{\theta}}(s,a)=A^{soft}_{f_{\theta}}(s,a)$, which also implies that $\pi^{*}_{f_{\theta}}=\pi^{*}_{R_{\theta}}$. Then, we can deduce the gradient of the loss of discriminator.
$-\nabla_{\theta}\mathcal{L}_{disc}$ $=\mathbb{E}_{\mathcal{D}_{exp}}\left[\nabla_{\theta}\log D_{\theta}(s,a)\right]+\mathbb{E}_{\pi}\left[\nabla_{\theta}\log(1-D_{\theta}(s,a))\right]$
$=\mathbb{E}_{\mathcal{D}_{exp}}\left[\frac{1}{2}\nabla_{\theta}f_{\theta}(s,a)\right]-\mathbb{E}_{\pi}\left[\frac{1}{2}\nabla_{\theta}f_{\theta}(s,a)\right],$
$-2\nabla_{\theta}\mathcal{L}_{disc}$ $=\mathbb{E}_{\mathcal{D}_{exp}}\left[\nabla_{\theta}f_{\theta}(s,a)\right]-\mathbb{E}_{\pi}\left[\nabla_{\theta}f_{\theta}(s,a)\right].$
Appendix C Performance Gap Analysis
Lemma C.1 (Implicit Feasible Reward Set (Ng et al., 2000)).

### v2

@@ 6 Conclusion
We thank a bunch of people.
Appendix A Related Works
Generative IRL.
Margin optimization based IRL methods (ng2000algorithms; abbeel2004apprenticeship; ratliff2006maximum) aim to learn reward functions that explain expert behavior better than other policies by a margin. Bayesian approaches were introduced with different prior assumptions on reward distributions, such as Boltzmann distributions (ramachandran2007bayesian; choi2011map; chan2021scalable) or Gaussian Processes (levine2011nonlinear). To avoid biases from maximum margin methods, (ziebart2008maximum; ziebart2010modeling) proposed a Lagrangian dual framework to cast the reward learning into a maximum likelihood problem with linear-weighted feature-based reward representation. wulfmeier2015maximum extended the framework to nonlinear reward representations, and finn2016guided combined importance sampling techniques to enable model-free estimation. Inspired by GANs, generative methods were introduced for policy and reward learning in IRL (ho2016generative; fu2017learning; torabi2018generative). However, these methods typically work with Maximum Entropy (ME) formulation yet suffer from sample inefficiency and stochasticity. Although there have been efforts to combine generative methods with off-policy RL agents to improve sample efficiency (kostrikov2018discriminator; blonde2019sample; blonde2022lipschitzness), few extend it to the model-based setting which might further the improvement, and none of these approaches addresses the rewards learning in stochastic MDP.
MBIRL.
Integrating IRL with MBRL has also shown success. For example, das2021model and herman2016inverse presented a gradient-based IRL approach using different policy optimization methods with dynamic models for linear-weighted features reward learning. In das2021model, the dynamic model is used to pass forward/backward the gradient in order to update the IRL and policy optimization modules. Similarly, end-to-end differentiable adversarial IRL frameworks to various state spaces have also been explored (baram2016model; baram2017end; sun2021adversarial; rafailov2021visual), where dynamic model serves a similar role. Despite these advancements, existing methods rarely address the specific challenges posed by stochastic environments, which limit reward learning performance.
Reward Shaping.
Reward shaping (dorigo1994robot; randlov1998learning) is a technique that enhances the original reward signal by adding additional domain information, making it easier for the agent to learn optimal behavior. This can be defines as $\hat{R}=R+F$, where $F$ is the shaping function and $\hat{R}$ is the shaped reward function. Potential-based reward shaping (PBRS) (ng2000algorithms) builds the potential function on states, $F(s,a,s^{\prime})=\phi(s^{\prime})-\phi(s)$, while ensuring the policy invariance property, which refers to inducing the same optimal behavior under different rewards $R$ and $\hat{R}$. Nonetheless, there exists other variants on the inputs of the potential functions such as state-action (wiewiora2003principled), state-time (devlin2012dynamic), and value function (harutyunyan2015expressing) as potential function input. There are also some recent attempts of reward shaping without utilization of domain knowledge potential function to solve exploration under sparse rewards (hu2020learning; devidze2022exploration; gupta2022unpacking; skalse2023invariance).
Appendix B Reward Shaping Soft Optimal Policy
Theorem B.1.
@@ Appendix B Reward Shaping Soft Optimal Policy
Proof B.2.
According to Soft VI ( section 2), we can expand the representation of $Q^{soft}_{\hat{R}}(s_{t},a_{t})$ as follows.
@@
From above derivation, we can tell that $Q^{soft}_{\hat{R}}(s_{t},a_{t})+\phi(s_{t})$ satisfy the soft bellmen update with original $R$. Thus, with simple induction, we can arrive that $Q_{R}^{soft}(s_{t},a_{t})=Q_{\hat{R}}^{soft}(s_{t},a_{t})+\phi(s_{t})$. Then, we can derive the advantage function
@@
Appendix C Adversarial Reward Learning
Proposition C.1.
Consider an undiscounted MDP. Suppose $f_{\theta}$ and $\pi$ at current iteration are the soft-optimal advantage function and policy for reward function $R_{\theta}$. Minimising the cross-entropy loss of the discriminator under generator $\pi$ is equivalent to maximising the log-likelihood under Maximum Causal Entropy IRL.
Proof C.2.
@@ Appendix C Adversarial Reward Learning
$=\mathbb{E}_{\mathcal{D}_{exp}}\left[\sum_{t=0}^{T-1}\log\pi(a_{t}|s_{t})+\log\rho_{0}+\sum_{t=1}^{T}\log\mathcal{T}(s_{t+1}|s_{t},a_{t})\right]$
@@
Breaking down above equations with soft VI (section 2), we can arrive the following.
@@
In our case, the transition function $\mathcal{T}$ is estimated by an approximation function $\hat{\mathcal{T}}$, which is updated with samples from $\mathcal{D}_{exp}$ and samples from off-policy buffer $\mathcal{D}_{env}$, thus we can drop the $\mathbb{E}_{\mathcal{T}}$ here. And $B$ term will cancel out, ending up to 0. To summarize, the gradient of log MLE loss of MCE IRL is the following.
@@
$\log D_{\theta}(s,a,\mathcal{T})$ $=f_{\theta}(s,a,\mathcal{T})-\log(\exp\{f_{\theta}(s,a,\mathcal{T})\}+\pi(a|s)),$
$\log(1-D_{\theta}(s,a,\mathcal{T}))$ $=\log\pi(a|s)-\log(\exp\{f_{\theta}(s,a,\mathcal{T})\}+\pi(a|s)).$
@@
$\nabla_{\theta}\log D_{\theta}(s,a,\mathcal{T})$ $=\nabla_{\theta}f_{\theta}(s,a,\mathcal{T})-\frac{\exp\{f_{\theta}(s,a,\mathcal{T})\}\nabla_{\theta}f_{\theta}(s,a,\mathcal{T})}{\exp\{f_{\theta}(s,a,\mathcal{T})\}+\pi(a|s)},$
$\nabla_{\theta}\log(1-D_{\theta}(s,a,\mathcal{T}))$ $=-\frac{\exp\{f_{\theta}(s,a,\mathcal{T})\}\nabla_{\theta}f_{\theta}(s,a,\mathcal{T})}{\exp\{f_{\theta}(s,a,\mathcal{T})\}+\pi(a|s)}.$
Since $\pi$ is trained by using $f_{\theta}$ as shaped reward, from soft VI we can derive that $\pi^{*}_{f_{\theta}}(a|s)=\exp A^{soft}_{f_{\theta}}(s,a)$. By assumption, we assume that $f_{\theta}$ is the advantage function of $R_{\theta}$, $f_{\theta}(s,a)=A^{soft}_{R_{\theta}}(s,a)$. From Theorem 3.2, we know that $A^{soft}_{R_{\theta}}(s,a)=A^{soft}_{f_{\theta}}(s,a)$, which also implies that $\pi^{*}_{f_{\theta}}=\pi^{*}_{R_{\theta}}$. Then, we can deduce the gradient of the loss of discriminator.
$-\nabla_{\theta}\mathcal{L}_{disc}$ $=\mathbb{E}_{\mathcal{D}_{exp}}\left[\nabla_{\theta}\log D_{\theta}(s,a,\mathcal{T})\right]+\mathbb{E}_{\pi}\left[\nabla_{\theta}\log(1-D_{\theta}(s,a,\mathcal{T}))\right]$
$=\mathbb{E}_{\mathcal{D}_{exp}}\left[\frac{1}{2}\nabla_{\theta}f_{\theta}(s,a,\mathcal{T})\right]-\mathbb{E}_{\pi}\left[\frac{1}{2}\nabla_{\theta}f_{\theta}(s,a,\mathcal{T})\right],$
$-2\nabla_{\theta}\mathcal{L}_{disc}$ $=\mathbb{E}_{\mathcal{D}_{exp}}\left[\nabla_{\theta}f_{\theta}(s,a,\mathcal{T})\right]-\mathbb{E}_{\pi}\left[\nabla_{\theta}f_{\theta}(s,a,\mathcal{T})\right].$
Appendix D Performance Gap Analysis
Lemma D.1 (Implicit Feasible Reward Set (ng2000algorithms)).